OpenStudy (anonymous):

c squared = a squared + b squared with c=32 with the other sides length 4:3 what is length of the two sides

OpenStudy (kc_kennylau):



OpenStudy (jack1):

simultaneous equations so \[\large c^2 = a^2 + b^2 \] and \[\large a = \frac 43 b \]

OpenStudy (jack1):

\[\large 32^2 = a^2 + b^2 \] sub eqn 2 into eqn 1 \[\large 32^2 = (\frac43 b)^2 + b^2 \] can u solve from here? @firetiger ?

OpenStudy (kc_kennylau):

I would let \(a\) as \(3x\) and \(b\) as \(4x\) since it's much easier to comprehend :)

OpenStudy (jack1):

agreed @kc_kennylau @firetiger use ken's diagram, it's probably easier

OpenStudy (anonymous):

ok square root(32)=(4x)squared +(3x) squared is the correct equation?

OpenStudy (shamil98):

\[\large 32^2 = 3x + 4x\]

OpenStudy (kc_kennylau):

\[32^2=(4x)^2+(3x)^2\]

OpenStudy (shamil98):

oh yeah forgot bout the squared part

OpenStudy (kc_kennylau):

Sorry I got to go see you guys later (others will help him)

OpenStudy (anonymous):

Ok the background a tv measures 32" across the screen the width is 4x with the height is 3x. x=?

OpenStudy (shamil98):

1024 = 16x^2 + 9x^2 solve for x.

OpenStudy (jack1):

did u get an answer for x @firetiger ?